# Self-induced stochastic resonance: A physics-informed machine learning approach

### v1

@@ Abstract
Self-induced stochastic resonance (SISR) is the emergence of coherent oscillations in slow-fast excitable systems driven solely by noise, without external periodic forcing or proximity to a bifurcation. This work presents a physics-informed machine learning framework for modeling and predicting SISR in the stochastic FitzHugh–Nagumo neuron. We embed the governing stochastic differential equations and SISR-asymptotic timescale-matching constraints directly into a Physics-Informed Neural Network (PINN) based on a Noise-Augmented State Predictor architecture. The composite loss integrates data fidelity, dynamical residuals, and barrier-based physical constraints derived from Kramers’ escape theory. The trained PINN accurately predicts the dependence of spike-train coherence on noise intensity, excitability, and timescale separation, matching results from direct stochastic simulations with substantial improvements in accuracy and generalization compared with purely data-driven methods, while requiring significantly less computation. The framework provides a data-efficient and interpretable surrogate model for simulating and analyzing noise-induced coherence in multiscale stochastic systems.
@@ 1 Introduction
Noise, often regarded as a source of disorder and unpredictability, can in fact play a constructive role in nonlinear dynamical systems [1, 2]. In a variety of physical, biological, and engineering contexts, random fluctuations can induce or enhance regularity in systems that are otherwise quiescent or weakly active. This counterintuitive phenomenon—where noise promotes order rather than destroys it—has been extensively studied under the umbrella of stochastic resonance and its generalizations [3, 4, 5, 6, 7]. In excitable systems such as neurons, noise can trigger coherent oscillations that would not occur in the absence of random fluctuations, thereby shaping information transmission and signal reliability in the nervous system [8].
@@
Among the diverse manifestations of noise-induced coherence, self-induced stochastic resonance (SISR) [9, 10, 11, 12, 13, 14] occupies a particularly distinctive position. Classical stochastic resonance (SR) requires the simultaneous presence of noise and a weak periodic forcing [3, 4], with the noise optimally amplifying the external signal. Coherence resonance (CR) [15, 16, 17], in contrast, arises in autonomous systems tuned close to a bifurcation threshold [18], where noise enhances the intrinsic tendency toward oscillations without any external drive. SISR differs fundamentally from both: it occurs in excitable systems whose deterministic dynamics possess only stable fixed point(s), far from any bifurcation [11, 17], no external periodic forcing is required, and, unlike SR and CR, it requires a strong timescale separation between the system variables. In this regime, no deterministic oscillations exist, yet weak stochastic perturbations alone can generate highly regular, self-sustained oscillations.
@@
The mechanism behind the SISR phenomenon lies in the matching between two characteristic timescales—the deterministic relaxation along the slow manifold and the stochastic escape time across potential barriers—first elucidated through Kramers’ escape-rate theory and large deviation analysis [11, 12]. When these timescales coincide, noise-induced transitions occur at reproducible points in phase space, producing coherent rhythmic activity. Thus, SISR reveals how random fluctuations, rather than an external drive or proximity to instability, can organize slow–fast dynamics into emergent periodic behavior, bridging nonequilibrium statistical physics and neuronal excitability [12, 13, 14, 19, 20, 21].
@@
Despite its conceptual significance, SISR remains analytically and numerically challenging to characterize. The phenomenon is inherently multiscale and stochastic: accurate prediction of coherence requires resolving both the slow deterministic drift of the recovery variable and the exponentially rare noise-induced escapes from potential wells. Classical numerical approaches, such as direct integration of the stochastic FitzHugh–Nagumo (FHN) equations, can reproduce SISR but at the cost of long simulations and extensive ensemble averaging to capture rare transitions [12, 20, 22, 23]. Analytical approaches, including asymptotic analysis [1, 24], large-deviation theory [9, 10, 25], and Kramers’ law [26], provide deep theoretical insight into escape dynamics and resonance conditions, but their validity is often restricted to idealized limits (e.g., weak noise, infinite timescale separation), limiting applicability to realistic parameter regimes.
@@
Recent advances in scientific machine learning have opened new avenues for addressing such multiscale stochastic problems [27, 28, 29, 30, 31]. In particular, physics-informed neural networks (PINNs) offer a promising framework by embedding governing stochastic multiple timescale equations and physical constraints directly into the learning process [32, 33, 34, 35]. By combining data-driven inference with dynamical priors derived from the system’s equations, PINNs can efficiently approximate solutions to stochastic differential equations even when only limited or noisy data are available. This hybrid approach enables not only the reconstruction of stochastic trajectories but also the identification of physically meaningful quantities such as potential barriers, escape rates, and timescale ratios that govern the emergence of SISR.
@@
This work presents, for the first time, a PINN framework for modeling and predicting SISR in an excitable FHN neuron. It is worth mentioning that the prediction of coherence resonance, a distinct noise-induced resonance phenomenon from SISR [17], has previously been addressed using a machine learning approach based on physics-unaware reservoir computing [36], a paradigm that differs fundamentally from PINNs. In contrast, the framework proposed in the present study incorporates the stochastic dynamics of the system directly into the training objective of the feedforward network, while simultaneously enforcing the physics constraints necessary for the emergence of SISR. Specifically, we design a composite loss function that combines data consistency, dynamical residuals, and constraints reflecting the asymptotic matching between deterministic and stochastic timescales as predicted by Kramers’ theory. Through this formulation, the neural network not only learns to reproduce the observed stochastic trajectories but also internalizes the physical structure that governs coherent noise-induced oscillations.
@@
Our results demonstrate that the physics-informed framework can accurately capture the onset and degree of SISR from simulated data while significantly reducing computational cost compared to traditional numerical simulations. Furthermore, by comparing different configurations of the loss functional, we show that embedding physical knowledge directly into the learning process enhances both interpretability and generalization. The proposed method thus establishes a general paradigm for integrating stochastic dynamical theory with neural network learning—extending the reach of PINNs to complex noise-driven phenomena such as self-induced stochastic resonance.
@@ 3 SISR: The physics approach
In this section, we analyze SISR from the perspective of nonlinear dynamics, viewing it as a noise-driven phenomenon emerging from the interplay between fast–slow dynamics and stochastic perturbations in excitable systems. We characterize the potential landscape, energy barriers, and noise-induced transitions that underpin the coherent spiking behavior of SISR. This physics-based understanding provides the theoretical foundation for the physics-informed framework developed in the next section.
We consider the stochastic FHN neuron of Eq. (1) in the excitable regime $0<a<1+\sqrt{2}$, where the deterministic dynamics admit only a stable fixed point. Our goal is to examine how stochastic forcing of amplitude $\sigma$ generates coherent spikes in the absence of any deterministic oscillations.
@@
The stationary probability flux $j_{0}$ in the limit $\sigma\to 0$ yields the escape rate in Arrhenius form [1, 43, 12]:
@@
In Eq. (19) The first condition guarantees uniqueness and stability of the fixed point; the second ensures that escape from the left stable branch of the $v$-nullcline occurs at $w=w_{\ell}$ before trajectories are trapped in the basin of attraction of the fixed point $(v_{e},w_{e})$; the third ensures almost-sure escape from the left (right) branch of the $v$-nullcline at $w=w_{\ell}$ ($w_{r}$); and the fourth ensures uniqueness of $w_{\ell,r}$ through the monotonicity of $\Delta U_{\ell,r}(w,a)$ (Fig. 3), leading to periodic, coherent spiking. The coherence of the spike train depends on how accurately the third condition is fulfilled, distinguishing SISR from coherence resonance, which instead requires parameter tuning near a bifurcation [17, 36].
@@ 4 Numerical results: effect of excitability and timescale separation on SISR
A value $\mathrm{CV}=1$ indicates a Poisson-like (incoherent) spike train, while $\mathrm{CV}>1$ corresponds to super-Poissonian irregularity [47]. In both cases, the system is far from the synchronization condition of Eq. (19). Conversely, as $\mathrm{CV}\to 0$, the ISIs become highly regular and the SISR condition is increasingly well satisfied. Perfectly periodic spiking corresponds to $\mathrm{CV}=0$, while small but nonzero values of $\mathrm{CV}$ quantify strong SISR.
@@
Figure 4 shows the influence of $a\in(0,1+\sqrt{2})$ on SISR at fixed $\varepsilon=0.00025$. Increasing $a$ reduces both the frequency and regularity of noise-induced spikes, as evident in the time series (a1)–(a3) and phase portraits (b1)–(b3). The fixed point, given by the intersection of the $v$- and $w$-nullclines, moves farther from the minimum of the $v$-nullcline as $a$ increases, enhancing excitability (i.e., rendering the stable fixed point more strongly attracting) but degrading temporal coherence. This behavior is quantified in Fig. 4(c): smaller $a$ yields lower $\mathrm{CV}$ minima at weaker noise, while larger $a$ shifts the minima to higher $\sigma$ values, consistent with a loss of timescale matching in Eq. (19).
@@
Figure 5 explores the effect of $\varepsilon$ at fixed $a=0.05$. Increasing $\varepsilon$ leads to more frequent but less coherent spiking, without changing the fixed point location, as seen in panels (b1)–(b3). This reflects a purely dynamical effect: larger $\varepsilon$ shortens the deterministic timescale $\varepsilon^{-1}$ along the stable nullcline branches, thereby disrupting the matching condition in Eq. (19). As shown in Fig. 5(c), $\mathrm{CV}$ minima occur at smaller $\varepsilon$, while larger $\varepsilon$ shifts the minima toward higher noise amplitudes and increases overall $\mathrm{CV}$, confirming the degradation of coherence.
@@
While these simulations validate the theoretical SISR conditions, direct numerical integration is computationally expensive, requiring long simulations to resolve rare stochastic transitions. Moreover, such methods struggle with noisy, sparse, or incomplete data, which is typical of experimental settings. PINNs provide an efficient alternative by directly embedding the governing equations into the learning objective, enabling the data-driven reconstruction of SISR dynamics with limited data and under physical constraints. The next section develops this framework in detail.
@@ 5.2 Physics-informed neural networks
While conventional neural networks can approximate arbitrary nonlinear mappings, their accuracy depends heavily on large amounts of labeled data—often impractical in scientific contexts. PINNs overcome this limitation by embedding the governing differential equations directly into the loss functional [35]. This strategy combines data-driven learning with physics-based regularization, enabling efficient inference from sparse or noisy observations while maintaining consistency with the underlying system dynamics. In this sense, PINNs provide a principled bridge between stochastic dynamical modeling and data-driven learning, allowing the SISR mechanism to be reconstructed and analyzed within a unified framework.
@@ 5.3 Training data preparation
A defining feature of SISR is its emergence in the singularly perturbed limit $\varepsilon\to 0$, where the deterministic timescale $\varepsilon^{-1}$ becomes large. This slow relaxation along the nullcline is essential for timescale matching with the stochastic escape dynamics. However, this also poses a practical difficulty: accurately characterizing SISR requires a sufficiently large ensemble of interspike intervals (ISIs), which in turn demands exceedingly long numerical integrations to ensure statistical reliability in estimating spike-train coherence through measures such as the coefficient of variation ($\mathrm{CV}$).
This computational bottleneck motivates the use of PINNs, which can infer long-term stochastic dynamics from limited simulated data. Instead of performing prohibitively long simulations, we integrate the stochastic FHN system over a relatively short time horizon and use the resulting trajectories to train the network. Once trained, the PINN generalizes these short-time dynamics to predict the long-time statistics of the ISIs and thus the coefficient of variation ($\mathrm{CV}$), providing an efficient and physics-consistent estimate of the SISR coherence.
The training dataset consists of simulated trajectories $(v_{t},w_{t},\eta_{t})$ generated by numerically integrating Eq. (1) over the time interval $t\in[0,1.0\times 10^{4}]$ with time step $\Delta t=0.05$, yielding $N_{d}=2.0\times 10^{5}$ data points. Unless stated otherwise, the parameters are $a=0.05$, $b=1.0$, $c=2.0$, $\varepsilon=0.00025$, and $\sigma=0.03061$.

### v2

@@ Abstract
Self-induced stochastic resonance (SISR) is the emergence of coherent oscillations in slow-fast excitable systems driven solely by noise, without external periodic forcing or proximity to a bifurcation. This work presents a physics-informed machine learning framework for modeling and predicting SISR in the stochastic FitzHugh–Nagumo neuron. We embed the governing stochastic differential equations and SISR-asymptotic timescale-matching constraints directly into a Physics-Informed Neural Network (PINN) based on a Noise-Augmented State Predictor architecture. The composite loss integrates data fidelity, dynamical residuals, and barrier-based physical constraints derived from Kramers’ escape theory. The trained PINN accurately predicts the dependence of spike-train coherence on noise intensity, excitability, and timescale separation, matching results from direct stochastic simulations with substantial improvements in accuracy and generalization compared with purely data-driven methods, requiring significantly less computation. The framework provides a data-efficient and interpretable surrogate model for simulating and analyzing noise-induced coherence in multiscale stochastic systems.
@@ 1 Introduction
Noise, often regarded as a source of disorder and unpredictability, can in fact play a constructive role in nonlinear dynamical systems [1, 2, 3]. In a variety of physical, biological, and engineering contexts, random fluctuations can induce or enhance regularity in systems that are otherwise quiescent or weakly active. This counterintuitive phenomenon—where noise promotes order rather than destroys it—has been extensively studied under the umbrella of stochastic resonance and its generalizations [4, 5, 6]. In excitable systems such as neurons, noise can trigger coherent oscillations that would not occur in the absence of random fluctuations, thereby shaping information transmission and signal reliability in the nervous system [7].
@@
Among the diverse manifestations of noise-induced coherence, self-induced stochastic resonance (SISR) [8, 9, 10] occupies a particularly distinctive position. Classical stochastic resonance (SR) requires the simultaneous presence of noise and a weak periodic forcing [4, 11], with the noise optimally amplifying the external signal. Coherence resonance (CR) [12, 13, 14], in contrast, arises in autonomous systems tuned close to a bifurcation threshold [15], where noise enhances the intrinsic tendency toward oscillations without any external drive. SISR differs fundamentally from both: it occurs in excitable systems whose deterministic dynamics possess only stable fixed point(s), far from any bifurcation [8, 14], no external periodic forcing is required, and, unlike SR and CR, it requires a strong timescale separation between the system variables. In this regime, no deterministic oscillations exist, yet weak stochastic perturbations alone can generate highly regular, self-sustained oscillations.
@@
The mechanism behind the SISR phenomenon lies in the matching between two characteristic timescales—the deterministic relaxation along the slow manifold and the stochastic escape time across potential barriers—first elucidated through Kramers’ escape-rate theory and large deviation analysis [8, 9, 16]. When these timescales coincide, noise-induced transitions occur at reproducible points in phase space, producing coherent rhythmic activity. Thus, SISR reveals how random fluctuations, rather than an external drive or proximity to instability, can organize slow–fast dynamics into emergent periodic behavior, bridging nonequilibrium statistical physics and neuronal excitability [10, 17, 18].
@@
Despite its conceptual significance, SISR remains analytically and numerically challenging to characterize. The phenomenon is inherently multiscale and stochastic: accurate prediction of coherence requires resolving both the slow deterministic drift of the recovery variable and the exponentially rare noise-induced escapes from potential wells. Classical numerical approaches, such as direct integration of the stochastic FitzHugh–Nagumo (FHN) equations, can reproduce SISR but at the cost of long simulations and extensive ensemble averaging to capture rare transitions [16, 19, 20]. Analytical approaches, including asymptotic analysis [2, 21], large-deviation theory [1, 22, 23], and Kramers’ law [24], provide deep theoretical insight into escape dynamics and resonance conditions, but their validity is often restricted to idealized limits (e.g., weak noise, infinite timescale separation), limiting applicability to realistic parameter regimes.
@@
Recent advances in scientific machine learning have opened new avenues for addressing such multiscale stochastic problems [25, 26]. In particular, physics-informed neural networks (PINNs) offer a promising framework by embedding the governing stochastic multiple timescale equations and physical constraints directly into the learning process [27, 28, 29]. By combining data-driven inference with dynamical priors derived from the system’s equations, PINNs can efficiently approximate solutions to stochastic differential equations even when only limited or noisy data are available. This hybrid approach enables not only the reconstruction of stochastic trajectories but also the identification of physically meaningful quantities such as potential barriers, escape rates, and timescale ratios that govern the emergence of SISR.
Beyond PINNs, there is a rapidly growing body of work on machine-learning surrogates for stochastic and multiscale dynamical systems. Deep generative models and neural surrogates have been proposed to emulate stochastic simulators and reproduce output distributions, enabling fast uncertainty quantification and rare-event statistics [30]. For multiscale stochastic systems with explicit slow–fast structure, data-driven methods have been developed to learn effective dynamics for the slow variables [31], discover low-dimensional embeddings, and construct reduced models that retain the essential stochastic behavior of the full system, including escape events and regime transitions [32, 33]. In parallel, neural stochastic differential equations (neural SDEs) represent both drift and diffusion terms with neural networks and have emerged as flexible surrogates for complex stochastic processes in continuous time, combining the expressive power of deep networks with SDE-based priors on temporal evolution [34, 35]. Our NASP–PINN framework is complementary to these approaches: it focuses on excitable slow–fast neuron dynamics, explicitly encodes the SISR timescale-matching conditions via the barrier-based constraints , and leverages physics-informed training to obtain a noise-aware surrogate that faithfully reproduces coherent spiking induced solely by stochastic forcing.
@@
It is worth mentioning that the prediction of coherence resonance, a distinct noise-induced resonance phenomenon from SISR [14], has previously been addressed using a machine learning approach based on physics-unaware reservoir computing [36], a paradigm fundamentally different from PINNs. In contrast to previous studies that rely purely on data-driven learning or deterministic PINNs, we introduce a Physics-Informed Neural Network based on a Noise-Augmented State Predictor (NASP) that explicitly incorporates stochastic forcing into the state-transition map. This formulation embeds both stochastic differential dynamics and asymptotic escape-time constraints from Kramers’ theory directly into the composite loss. The PINN therefore learns noise-aware slow–fast dynamics from relatively short trajectories while respecting the underlying physics of noise-induced escapes. To our knowledge, this work presents the first PINN framework capable of modeling and predicting SISR, demonstrated here in the stochastic FHN neuron, although the method is in principle applicable to a broad class of multiscale stochastic systems.
@@
Our results demonstrate that the physics-informed framework can accurately capture the onset and degree of SISR from simulated data while significantly reducing computational cost compared to traditional numerical simulations. Although the FHN model can be solved numerically via standard stochastic integrators, our aim is not mere simulation but to develop a data-driven surrogate capable of reproducing both short-term dynamics and long-term stochastic statistics from limited trajectory data. Furthermore, by comparing different configurations of the loss functional, we show that embedding physical knowledge directly into the learning process enhances both interpretability and generalization. The proposed method thus establishes a general paradigm for integrating stochastic dynamical theory with neural network learning—extending the reach of PINNs to complex noise-driven phenomena such as self-induced stochastic resonance.
@@ 3 SISR: The physics approach
In this section, we analyze SISR from the perspective of nonlinear dynamics, viewing it as a noise-driven phenomenon emerging from the interplay between fast–slow dynamics and stochastic perturbations in excitable systems. We characterize the potential landscape, energy barriers, and noise-induced transitions that underpin the coherent spiking behavior of SISR. This physics-based understanding provides the theoretical foundation for the physics-informed neural network framework developed in the next section.
We consider the stochastic FHN neuron of Eq. (1) in the excitable regime, $0<a<1+\sqrt{2}$, where the deterministic dynamics admit only a stable fixed point. Our goal is to examine how stochastic forcing of amplitude $\sigma$ generates coherent spikes in the absence of any deterministic oscillations.
@@
The stationary probability flux $j_{0}$ in the limit $\sigma\to 0$ yields the escape rate in Arrhenius form [2, 43]:
@@
In Eq. (19), the first condition guarantees uniqueness and stability of the fixed point; the second ensures that escape from the left stable branch of the $v$-nullcline occurs at $w=w_{\ell}$ before trajectories are trapped in the basin of attraction of the fixed point $(v_{e},w_{e})$; the third ensures almost-sure escape from the left (right) branch of the $v$-nullcline at $w=w_{\ell}$ ($w_{r}$); and the fourth ensures uniqueness of $w_{\ell,r}$ through the monotonicity of $\Delta U_{\ell,r}(w,a)$ (Fig. 3), leading to periodic, coherent spiking. The coherence of the spike train depends on how accurately the third condition is fulfilled, distinguishing SISR from coherence resonance, which instead requires parameter tuning near a bifurcation [14, 36].
@@ 4 Numerical results: effect of excitability and timescale separation on SISR
A value $\mathrm{CV}=1$ indicates a Poisson-like (incoherent) spike train, whereas $\mathrm{CV}>1$ corresponds to super-Poissonian irregularity [47]. In both cases, the system is far from the synchronization condition of Eq. (19). Conversely, as $\mathrm{CV}\to 0$, the ISIs become highly regular and the SISR condition is increasingly well satisfied. Perfectly periodic spiking corresponds to $\mathrm{CV}=0$, and small but nonzero values of $\mathrm{CV}$ quantify strong SISR. Furthermore, we note that $\mathrm{CV}$ becomes mathematically meaningful (i.e., informative about spike-train coherence) only when at least three ISIs are available; however, a reliable estimate typically requires many more ISIs (on the order of $\sim 10$ for a coarse estimate).
@@
Figure 4 shows the influence of $a\in(0,1+\sqrt{2})$ on SISR at fixed $\varepsilon=0.00025$. Increasing $a$ reduces both the frequency and regularity of noise-induced spikes, as evident in the time series (a1)–(a3) and phase portraits (b1)–(b3). The fixed point, given by the intersection of the $v$- and $w$-nullclines, moves farther from the minimum of the $v$-nullcline as $a$ increases, enhancing excitability (i.e., rendering the stable fixed point more strongly attracting) but degrading temporal coherence. This behavior is quantified in Fig. 4(c): smaller $a$ yields lower $\mathrm{CV}$ minima at weaker noise, whereas larger $a$ shifts the minima to higher $\sigma$ values, consistent with a loss of timescale matching in Eq. (19).
@@
Figure 5 explores the effect of $\varepsilon$ at fixed $a=0.05$. Increasing $\varepsilon$ leads to more frequent but less coherent spiking, without changing the fixed point location, as seen in panels (b1)–(b3). This reflects a purely dynamical effect: larger $\varepsilon$ shortens the deterministic timescale $\varepsilon^{-1}$ along the stable nullcline branches, thereby disrupting the matching condition in Eq. (19). As shown in Fig. 5(c), $\mathrm{CV}$ minima occur at smaller $\varepsilon$, whereas larger $\varepsilon$ shifts the minima toward higher noise amplitudes and increases overall $\mathrm{CV}$, confirming the degradation of coherence.
@@
Although these simulations validate the theoretical SISR conditions, direct numerical integration is computationally expensive, requiring long simulations to resolve rare stochastic transitions. Moreover, such methods struggle with noisy, sparse, or incomplete data, which is typical of experimental settings. PINNs provide an efficient alternative by directly embedding the governing equations into the learning objective, enabling the data-driven reconstruction of SISR dynamics with limited data and under physical constraints. The next section develops this framework in detail.
@@ 5.2 Physics-informed neural networks
Although conventional neural networks can approximate arbitrary nonlinear mappings, their accuracy depends heavily on large amounts of labeled data—often impractical in scientific contexts. PINNs overcome this limitation by embedding the governing differential equations directly into the loss functional [54]. This strategy combines data-driven learning with physics-based regularization, enabling efficient inference from sparse or noisy observations while maintaining consistency with the underlying system dynamics. In this sense, PINNs provide a principled bridge between stochastic dynamical modeling and data-driven learning, allowing the SISR mechanism to be reconstructed and analyzed within a unified framework.
@@
It is worth noting that our goal is not merely to numerically solve the stochastic FHN model—which can already be done using a standard Euler-Maruyama integrator—but rather to learn a data-driven surrogate that reproduces both short-term dynamics and long-term stochastic statistics of SISR from relatively short simulated trajectories. The NASP–PINN learns the stochastic one-step transition map $(v_{t},w_{t},\sigma\eta_{t})\mapsto(v_{t+\Delta t},w_{t+\Delta t})$ and can subsequently generate long rollouts, ISI statistics, and $\mathrm{CV}$-noise curves without reintegrating the SDE, acting as a learned Euler-Maruyama-type integrator with reduced computational cost compared to Monte Carlo simulation.
The physics losses serve as inductive priors rather than equation replacements: phy1 enforces consistency with the deterministic vector field, while phy2 implements the Kramers timescale-matching mechanism underlying SISR, jointly improving training and generalization. This work, therefore, operates in the physics-known regime, where the full FHN equations provide a controlled environment to demonstrate the benefit of physically guided learning. In more realistic scenarios with unknown or incomplete physical knowledge (e.g., experimental recordings), phy1/phy2 could be replaced or complemented by alternative priors such as data-driven drift/diffusion estimation, nullcline inference, escape-time constraints, SINDy-based sparse regression [57, 58], or other structural properties. This study thus establishes a benchmark for physics-regularized surrogate modeling of SISR and motivates future extensions to partially known systems.
@@ 5.3 Training data preparation
A defining feature of SISR is its emergence in the singularly perturbed limit $\varepsilon\to 0$, where the deterministic relaxation timescale scales as $\varepsilon^{-1}$ and becomes large. The resulting slow drift along the vv–nullcline enables timescale matching with stochastic escapes, but also introduces a practical challenge: accurately estimating spike-train coherence (e.g., via the coefficient of variation, $\mathrm{CV}$) requires a sufficiently large ensemble of interspike intervals (ISIs), which demands long numerical integrations. As discussed earlier, $\mathrm{CV}$ only becomes meaningful once at least $3$ ISIs are available; however, reliable estimation generally requires significantly more samples ($\geq 10$ ISIs depending on variability). In practice, collecting even 551010 ISIs in the adiabatic regime $\varepsilon\to 0$ may require integration over $(15\times 10^{3}$$25\times 10^{3})$ time units, since spike events become increasingly rare as $\varepsilon$ decreases.
This computational bottleneck motivates the use of our NASP–PINN. Rather than performing prohibitively long simulations, we train the network using comparatively short trajectories of $T=10\times 10^{3}$ time units ($2\times 10^{5}$ steps with $dt=0.05$), which contain only $3$ ISIs for $\varepsilon=0.00025$. Such a window is insufficient for a reliable $\mathrm{CV}$ estimate, but provides enough local dynamical variability for NASP–PINN to learn the stochastic transition law. Once trained, the model is rolled out autonomously to generate long-horizon trajectories from which reliable $\mathrm{CV}$ can be estimated efficiently and without further numerical integration.
Unless stated otherwise, the training data consist of simulated trajectories $(v_{t},w_{t},\eta_{t})$ generated by numerically integrating Eq. (1) over $t\in[0,1.0\times 10^{4}]$ with time step $\Delta t=0.05$, yielding $N_{d}=2\times 10^{5}$ data points. Parameters are $a=0.05$, $b=1.0$, $c=2.0$, $\varepsilon=0.00025$, and $\sigma=0.03061$.